It takes Gary 1 hr. to milk all the cows, and it takes Dana 1.5 hrs. How long will it take them to do the job together?
Answer by richwmiller(17219)

3x/3+2x/3=1
3x+2x=3
5x=3
x=3/5*60=36 minutes